OpenStudy (anonymous):

What is the equation of the line passing through the points (3, 2) and (3, -4) ? I think it's undefined...but apparently I was incorrect, the answer is 3. Someone help please? What am I missing here?

OpenStudy (anonymous):

m = (y2 - y1) / (x2 - x1) m = (-4-2)/(3-3) = -6/0 undefined The slope is undefined... which means that this forms a vertical line (graph it). Vertical lines have the form x = some number... so answer here x = 3 is the equation of this particular vertical line.

OpenStudy (anonymous):

thanks!! for some reason i thought it said slope!! that's why i got it wrong -______-